# Overlapping signals in next-generation gravitational wave observatories: A recipe for selecting the best parameter estimation technique

### v2

@@ II Parameter estimation methods
[…]
@@ II.1 Single-signal parameter estimation
The simplest way to handle the actual PE of overlapping signals is to ignore the presence of a second signal, assuming a template for a unique merger as in Eq. (5). Generally, we recover only the parameters of the loudest signal present in the data at that time, possibly with biased estimates due to the other underlying signals. These biases can potentially be sufficiently large that the true parameters are outside of the recovered distribution. If the signals are close in SNR, we may recover the parameters of the quieter signal instead. For injections performed in this paper, we have never observed the sampler mxing both signals in the posterior.
@@ II.3 Joint parameter estimation
Of note is that the above likelihood is symmetric under the exchange of signals $A$ and $B$, so the posterior will inherit this symmetry. Still, for PE it is useful to break this symmetry between signals completely, to accurately recover the properties of each signal. To this end, we sort the posteriors to enforce an ordering of chirp masses to follow the ordering in the injection. For example $\mathcal{M}_{c,A,i}>\mathcal{M}_{c,B,i}$ for each sample $i$. This works for signals considered in this work, but for signals with very close chirp masses, some other method of ordering the parameters would need to be employed.
@@ II.4 Relative binning
$+B^{1}_{(l,m),(\bar{l},\bar{m})}(b)\left[s^{0}_{(l,m)}s^{0*}_{(\bar{l},\bar{m})}\left(r^{0}_{(l,m)}r^{1*}_{(\bar{l},\bar{m})}+r^{0*}_{(\bar{l},\bar{m})}r^{1}_{(l,m)}\right)\right.$
$+\left.\left.r^{0}_{(l,m)}r^{0*}_{(\bar{l},\bar{m})}\left(s^{0}_{(l,m)}s^{1*}_{(\bar{l},\bar{m})}+s^{0*}_{(\bar{l},\bar{m})}s^{1}_{(l,m)}\right)\right]\right\},$ (16)
@@ II.6 Tests of joint parameter estimation with relative binning
For our tests of JPE with RB, we use a detector network comprised of one 10 km triangular ET observatory located at the Virgo site and two CEs located at sites of LIGO-Hanford and LIGO-Livingston having arm lengths of 40 km and 20 km respectively. We inject two BBH signals with the IMRPhenomXPHM waveform model Pratten et al. (2021) into Gaussian noise following projected noise curves for ET Hild et al. (2011) and CE Abbott et al. (2017). We set the lower frequency cutoff of the analysis at 5 Hz, for which we note that traditional PE approaches become untractable for large injection studies. We then analyze the signals with the Bilby Ashton et al. (2019); Romero-Shaw et al. (2020) software using Dynesty Speagle (2019) as nested sampler. We use our custom implementation of the RB likelihood  [^1] https://github.com/corvus314/jointRB.
@@
Table 1: Summary of generation and analysis priors for tests of JPE. Both overlapping signals share the same priors, with signal $A$ chosen to have higher SNR in the pair. The parameters have their usual meanings: chirp mass $\mathcal{M}_{c}$, mass ratio $q$, dimensionless spins $a_{i}$ together with their angles $\theta_{i}$ to the orbital angular momentum, azimuthal angle between them $\Delta\phi$ and the precession angle $\phi_{JL}$. RA and DEC stand for the right ascension and the declination while $D_{L}$ is the luminosity distance to the source. $\theta_{JN}$ is the angle between the line of sight and the total angular momentum while $\psi$ is the polarization angle, while $\phi_{c}$ and $t_{c}$ stand for the phase and time at coalescence. The UniformInComponents prior on $\mathcal{M}_{c}$ and $q$ is equivalent to uniform priors in the component masses $m_{1},\,m_{2}$ within the prior boundaries. UniformSourceFrame is the distance prior uniform in comoving volume and the source frame time.
@@ II.6.1 Likelihood approximation
We start with assessing how well our RB JPE implementation approximates the true likelihood with the two signals in the strain. For constructing RB frequency bands, we follow Narola et al. (2023) for the reference waveform and bin settings, using the injected signal for the former, and a change in 2% for the chirp mass and mass ratio for the latter. We then recursively bisect the full frequency range until the RB error on the likelihood evaluated for the test waveform falls under a predetermined threshold, which we set to 0.03. These settings lead to a frequency grid of around a few hundred points, depending on the injected signal. We note that for some injections (around 1 in 5) this leads to a frequency grid that does not capture the likelihood properly, but allowing for a different percentage difference (typically 10%) in chirp mass and mass ratio fixes the problem.
@@ II.6.2 Performance on test injections
While this could suggest the less computationally expensive hierarchical subtraction could be used over JPE to get the posteriors quicker, we have to be careful as it does not perform well on all the events, like in the example posterior shown in Fig. 4. Here, when we recover overlapping signals with SSPE, the bias is more severe. While hierarchical subtraction reduces it substantially and shifts the posterior closer to JPE, it is not enough and the true values are still completely missed in the recovery. In principle, we could try more hierarchical subtraction steps to further improve the method, but more steps do not make sense from a computational point of view for two signals. With RB, JPE takes on average 11 days [^3] For our set of injections the average SNR was 53. on a 16-core CPU Intel Skylake Gold 6148, 6 times longer than a SSPE run and 2 times as long as the three PE runs needed for hierarchical subtraction. With two additional steps, most of the computational advantage of hierarchical subtraction would be lost. The method may still be useful where there are more than two signals overlapping in the strain.
@@ II.6.3 Percentile-percentile test
At the bottom, we show the result for recovery with JPE, where our data template properly models both mergers present in the strain. As we recover both signals, we plot 160 signals total. We see that the fraction of events in which we completely missed the injection in the recovery properly goes to 0. The p-values are much higher than for the previous methods with the lowest p-value again obtained for $a_{1}$ at 0.085. The combined p-value is 0.313, indicating that posteriors obtained from JPE are self-consistent.
@@ III.1 Prior-informed Fisher matrices
Fisher-based approaches are common for studies into the effects of overlapping signals in the 3G era Antonelli et al. (2021); Johnson et al. (2024). We incorporate the prior information into them and compare how well their predictions hold up agianst the PE posteriors.
@@ III.1.1 Test of the Fisher matrix method
For $\Sigma\lesssim 1$, so for the bias at or below the effects of the detector noise, the Fisher matrices fail to accurately predict the bias. The Fisher method can predict a relatively substantial bias of $\Sigma\approx 1$ even for small true bias in the $0.001-0.1$ range. As such it is not accurate enough to serve as a predictor of whether we need to perform JPE or just a standard PE. The standard deviation of the mean scales as $\sigma/\sqrt{N}$ where $\sigma$ is the standard deviation of the distribution and $N$ is the number of drawn samples. With $N$ in the $2000-10000$ range, the uncertainty in our definition of bias due to sampling is $\sim 0.01-0.02$. We would need the Fisher approach to be two orders of magnitude more accurate to predict small biases. These issues persist even when we expand near the true maximum likelihood point so we find Fisher formalism only useful when it predicts biases above the noise level $\Sigma\approx 1$.
@@ III.2 Time-frequency overlap
is the relative entropy Kullback and Leibler (1951) between the distributions. In  Romero-Shaw et al. (2020) the authors have shown that when sampling with Bilby package statistical fluctuations of the sampler lead to JSD of up to 0.0015 nats between the marginalized 1D posteriors. We adopt $\mbox{JSD}>0.0015$ as our criterion of bias due to presence of overlapping signals.
@@
We also plot the same results where we shifted the separation of the signals and kept all the other parameters the same, and we see an interesting pattern emerge. At $dt=0\,\mathrm{s}$ both signals are coincident at geocenter but it does not mean that they are coincident at the detectors. Depending on the sky positions of the signals, the ordering of the signals at the detectors can be different at each of the detectors. But as soon as the time difference between the mergers is larger than $0.045\,\mathrm{s}$ (the Earth’s diameter in seconds) the ordering of the signal at the interferometers is consistent between the interferometers and independent of sky location. The match and JSD are then driven primarily by ordering of the chirp masses of the signals, consistent with biases occurring due to overlaps in time-frequency. The chirp mass determines the leading order frequency evolution of the signal:
@@ III.3 Reweighting single-signal posteriors
If the reweighted posterior is the same as the original posterior (JSD $<0.0015$ nats for all 1D posteriors), the presence of overlapping signals was not important for inference. We can perform standard PE on signal B and we spend the same amount of computational resources as if there was no possibility of overlaps. If the posteriors are different, we run the full JPE analysis on the pair. Compared with starting from JPE from the beginning, computational cost increases around 17% by requiring the initial PE run. This does not pose a problem, as only a small fraction of mergers in the 3G era are expected to be affected by overlapping signals. It is also much lower than around 500% increase in the computational cost if we run JPE when only standard PE run is needed. In Fig. 11, we show a schematic flowchart of our proposed method.

### v3

@@ II Parameter estimation methods
with the inner product $\langle\cdot|\cdot\rangle$ defined as
@@ II.1 Single-signal parameter estimation
The simplest way to handle the actual PE of overlapping signals is to ignore the presence of a second signal, assuming a template for a unique merger as in Eq. (5). Generally, we recover only the parameters of the loudest signal present in the data at that time, possibly with biased estimates due to the other underlying signals. These biases can potentially be sufficiently large that the true parameters are outside of the recovered distribution. If the signals are close in SNR, we may recover the parameters of the quieter signal instead. For injections performed in this paper, we have never observed the sampler mixing both signals in the posterior. Sampling with Bilby uses differential evolution Ter Braak (2006) proposal method for drawing new nested sampling live points. Two overlapping signals have their modes far enough away in the parameter space that the sampler becomes fixated on one of the modes and does not propose jumps between them.
@@ II.3 Joint parameter estimation
Of note is that the above likelihood is symmetric under the exchange of signals $A$ and $B$, so the posterior will inherit this symmetry. Still, for PE it is useful to break this symmetry between signals completely, to accurately recover the properties of each signal. To this end, we sort the posteriors to enforce an ordering of chirp masses $\mathcal{M}_{c}$ to follow the ordering in the injection. We check if $\mathcal{M}_{\mathrm{inj},A}>\mathcal{M}_{\mathrm{inj},B}$, then for each sample $\theta_{i}$ we compute $(\mathcal{M}_{i,A}>\mathcal{M}_{i,B})=(\mathcal{M}_{\mathrm{inj},A}>\mathcal{M}_{\mathrm{inj},B})$. If true, we keep the sample as is; if false, we swap each parameter A with the corresponding parameter B for this sample. This works for signals considered in this work, but for signals with very close chirp masses, some other method of ordering the parameters would need to be employed.
@@ II.4 Relative binning
[…]
@@ II.6 Tests of joint parameter estimation with relative binning
For our tests of JPE with RB, we use a detector network comprised of one 10 km triangular ET observatory located at the Virgo site [^1] The design of ET has not been finalized between two L-shaped detectors and one triangular detector, the latter of which we chose for this work. For comparisons between designs, see Branchesi et al. (2023). and two CEs located at sites of LIGO-Hanford and LIGO-Livingston having arm lengths of 40 km and 20 km respectively. We inject two BBH signals with the IMRPhenomXPHM waveform model Pratten et al. (2021) into Gaussian noise following projected noise curves for ET Hild et al. (2011) and CE Abbott et al. (2017). We set the lower frequency cutoff of the analysis at 5 Hz, for which we note that traditional PE approaches become untractable for large injection studies. We then analyze the signals with the Bilby Ashton et al. (2019); Romero-Shaw et al. (2020) software using Dynesty Speagle (2019) as nested sampler. We use our custom implementation of the RB likelihood  [^2] https://github.com/corvus314/jointRB.
@@
Table 1: Summary of generation and analysis priors for tests of JPE. Both overlapping signals share the same priors, with signal AA chosen to have higher SNR in the pair. The parameters have their usual meanings: chirp mass $\mathcal{M}_{c}$, mass ratio $q$, dimensionless spins $a_{i}$ together with their angles $\theta_{i}$ to the orbital angular momentum, azimuthal angle between them $\Delta\phi$ and the precession angle $\phi_{JL}$. RA and DEC stand for the right ascension and the declination while $D_{L}$ is the luminosity distance to the source. $\theta_{JN}$ is the angle between the line of sight and the total angular momentum while $\psi$ is the polarization angle, while $\phi_{c}$ and $t_{c}$ stand for the phase and time at coalescence. The UniformInComponents prior on $\mathcal{M}_{c}$ and $q$ is equivalent to uniform priors in the component masses $m_{1},\,m_{2}$ within the prior boundaries. UniformSourceFrame is the distance prior uniform in comoving volume and the source frame time.
@@ II.6.1 Likelihood approximation
We start with assessing how well our RB JPE implementation approximates the true likelihood with the two signals in the strain. For constructing RB frequency bands, we follow Narola et al. (2023) for the reference waveform and bin settings, using the injected signal for the former, and a change in 2% for the chirp mass and mass ratio for the latter. We then recursively bisect the full frequency range until the RB error on the likelihood evaluated for the test waveform falls under a predetermined threshold, which we set to 0.03, following the default behavior of the RB code we adapted Janquart and Narola (2022) [^3] To be more precise, we require that the likelihood error in each bin falls below $0.03/\sqrt{N_{b}}$, where $N_{b}$ is the proposed number of bins at each iteration of the bisection algortihm. We run the algorithm until the number of bins and the requested number of bins converge. After our analysis, we discovered that the RB implementation we adapted sometimes stops this process while one more bisection is possible, due to a rounding issue. This has led to excess likelihood error over the desired level. Even with this bug, our likelihoods were accurate enough that none of the posteriors produced in this work were affected.. These settings lead to a frequency grid of around a few hundred points, depending on the injected signal. We note that for some injections (around 1 in 5) this leads to a frequency grid that does not capture the likelihood properly, but allowing for a different percentage difference (typically 10%) in chirp mass and mass ratio fixes the problem.
@@ II.6.2 Performance on test injections
While this could suggest the less computationally expensive hierarchical subtraction could be used over JPE to get the posteriors quicker, we have to be careful as it does not perform well on all the events, like in the example posterior shown in Fig. 4. Here, when we recover overlapping signals with SSPE, the bias is more severe. While hierarchical subtraction reduces it substantially and shifts the posterior closer to JPE, it is not enough and the true values are still completely missed in the recovery. In principle, we could try more hierarchical subtraction steps to further improve the method, but more steps do not make sense from a computational point of view for two signals. With RB, JPE takes on average 11 days [^5] For our set of injections the average SNR was 53. on a 16-core CPU Intel Skylake Gold 6148, 6 times longer than a SSPE run and 2 times as long as the three PE runs needed for hierarchical subtraction. With two additional steps, most of the computational advantage of hierarchical subtraction would be lost (although others have attempted to speed up HS steps using neural networks Hu (2025)). The method may still be useful where there are more than two signals overlapping in the strain.
@@ II.6.3 Percentile-percentile test
At the bottom, we show the result for recovery with JPE, where our data template properly models both mergers present in the strain. As we recover both signals, we plot 160 signals total. We see that the fraction of events in which we completely missed the injection in the recovery properly goes to 0. The p-values are much higher than for the previous methods with the lowest p-value again obtained for $a_{1}$ at 0.085. The combined p-value is 0.337, indicating that posteriors obtained from JPE are self-consistent.
@@ III.1 Prior-informed Fisher matrices
Fisher-based approaches are common for studies into the effects of overlapping signals in the 3G era Antonelli et al. (2021); Johnson et al. (2024). We incorporate the prior information into them and compare how well their predictions hold up against the PE posteriors.
@@ III.1.1 Test of the Fisher matrix method
For $\Sigma\lesssim 1$, so for the bias at or below the effects of the detector noise, the Fisher matrices fail to accurately predict the bias. The Fisher method can predict a relatively substantial bias of $\Sigma\approx 1$ even for small true bias in the 0.0010.10.001-0.1 range. As such it is not accurate enough to serve as a predictor of whether we need to perform JPE or just a standard PE. The standard deviation of the mean scales as $\sigma/\sqrt{N}$ where $\sigma$ is the standard deviation of the distribution and $N$ is the number of drawn samples. With $N$ in the 2000100002000-10000 range, the uncertainty in our definition of bias due to sampling is $\sim 0.01-0.02$. We would need the Fisher approach to be two orders of magnitude more accurate to predict small biases. These issues persist even when we expand near the true maximum likelihood point so we find Fisher formalism only useful when it predicts biases above the noise level $\Sigma\approx 1$.
@@ III.2 Time-frequency overlap
is the relative entropy Kullback and Leibler (1951) between the distributions. We chose JSD as our distance measure as it is symmetric between the distributions. In  Romero-Shaw et al. (2020) the authors have shown that when sampling with Bilby package statistical fluctuations of the sampler lead to JSD of up to 0.0015 nats between the marginalized 1D posteriors. We adopt $\mbox{JSD}>0.0015$ as our criterion of bias due to presence of overlapping signals.
@@
We also plot the same results where we shifted the separation of the signals and kept all the other parameters the same, and we see an interesting pattern emerge. At $dt=0\,\mathrm{s}$ both signals are coincident at geocenter but it does not mean that they are coincident at the detectors. Depending on the sky positions of the signals, the ordering of the signals at the detectors can be different at each of the detectors. But as soon as the time difference between the mergers is larger than $0.045\,\mathrm{s}$ (the Earth’s diameter in seconds) the ordering of the signal at the interferometers is consistent between the interferometers and independent of sky location. The match and JSD are then driven primarily by ordering of the chirp masses of the signals, consistent with biases occurring due to overlaps in time-frequency. The chirp mass determines the leading order frequency evolution of the signal Abbott et al. (2017):
@@ III.3 Reweighting single-signal posteriors
If the reweighted posterior is the same as the original posterior (JSD <0.0015<0.0015 nats for all 1D posteriors), the presence of overlapping signals was not important for inference. We can perform standard PE on signal B and we spend the same amount of computational resources as if there was no possibility of overlaps. If the posteriors are different, we run the full JPE analysis on the pair. Compared with starting from JPE from the beginning, computational cost increases around 17% by requiring the initial PE run. This does not pose a problem, as only a small fraction of mergers in the 3G era are expected to be affected by overlapping signals. It is also much lower than around 500% increase in the computational cost if we run JPE when only standard PE run is needed. In Fig. 11, we show a schematic flowchart of our proposed method.